OpenStudy (anonymous):

sovle the following equation for x x-2 /x+3 = x-2/x+6

OpenStudy (anonymous):

All you have to do for this is cross multiply, expand, and solve. So you have: 1) (x+6)*(x-2) = (x-2)*(x+3) 2) \[x ^{2}+4x-12 = x ^{2}+x-6\] 3) 3x-6 = 0 and so x = 2

OpenStudy (anonymous):

(x-2)(x+3)=(x-2)(x+6) =>x^2+x-6=x^2+4x-12 =>3x=6 =>x=2